OpenStudy (anonymous):

What is the directrix of the parabola given by the equation y = –5x2 + 60x – 176?

OpenStudy (anonymous):

answer choices: x= 3 19/20 y= 4 1/20 x=4 1/20 y= 2 19/20 Im completely lost on finding the directrix

OpenStudy (anonymous):

the last one is 3 19/20 (not 2 19/20)

OpenStudy (saifoo.khan):

Find the vertex, then find the focus. Finally directix.

OpenStudy (anonymous):

here we go again lets go slow

OpenStudy (anonymous):

please. I understand the focus more. But this confused me.

OpenStudy (anonymous):

\[y = -5x^2 + 60x -176\] what is your job? to make it look like \[4p(y-k)=(x-h)^2\] this is going to take some algebra

OpenStudy (anonymous):

first add \(176\) to both sides

OpenStudy (anonymous):

I just get confused with all the variables. I know how to solve equations and stuff but the formual confuses me

OpenStudy (anonymous):

actually i have a better idea to make it somewhat easier but you are going to have to complete the square no matter what

OpenStudy (anonymous):

okay thankyou!

OpenStudy (anonymous):

ok now i see i made a bush league mistake. lets start again

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[y = -5x^2 + 60x -176\] \[y=-5(x^2-12x)-176\] \[y=-5(x-6)^2+180-176\] \[y=-5(x-6)^2+4\] \[y-4=-5(x-6)^2\] \[-\frac{1}{5}(y-4)=(x-6)^2\]

OpenStudy (anonymous):

now it is in the form you want

OpenStudy (anonymous):

we see from our eyeballs that the vertex is \((6,4)\) and we compute \(p\) as before by solving \(4p=-\frac{1}{5}\) so \(p=-\frac{1}{20}\)

OpenStudy (anonymous):

okay I understand that. now how do you find the directrix though? Im not lost yet

OpenStudy (anonymous):

parabola opens down, so the focus is \(\frac{1}{20}\) units below the vertex and directrix is \(\frac{1}{20}\) units above

OpenStudy (anonymous):

focus is therefore \((6,4-\frac{1}{20})\) or \((6,3\tfrac{19}{20})\)

OpenStudy (anonymous):

and directrix is \(y=4+\frac{1}{20}\) or \(y=4\tfrac{1}{20}\)

OpenStudy (anonymous):

ohh okay! its starting to make more sense nnow! thankyou again for all your help!

OpenStudy (anonymous):



OpenStudy (anonymous):

you really need a picture to see what is going that is a crappy one here is a better one http://www.wolframalpha.com/input/?i=+y+%3D+%E2%80%935x2+%2B+60x+%E2%80%93+176

OpenStudy (anonymous):

if you take the link i sent, type "parabola" before the equation, you will get everything you want wont help on a test or quiz though